# Cubic polynomials and sums of two squares

### v1

@@ 5. Further Remarks
Theorem 1.1 concerns itself with $25\%$ of degree $3$ monic irreducible polynomials. This limitation arises because we only transferred lower bounds from $\mathcal{R}_{\theta,1}(\cdot,\cdot)$ to $B_{\mathbf{p}}(\cdot)$. It is important to note that $\mathcal{R}_{\theta,1}(\cdot,\cdot)$ can be empty, for example, if $a_{1}a_{2}-a_{0}$ is even. To cover all cases of irreducible monic polynomials of degree $3$, we expect one would need to use sets $\mathcal{R}_{\theta,2^{t}}(\cdot,\cdot)$ for a suitable integer $t$ and employ enhanced parity arguments; however, we do not pursue these additional complications here. Furthermore, our arguments should generalize to study integer points on the Châtelet surface $y^{2}+nz^{2}=\mathbf{p}(x)$. Here our theory is stronger when $n\in\mathbb{N}$, as our unit arguments can be transferred quite readily, but when $n$ is negative, most of our theory breaks down (with the exception of an analogue for Lemma 1.3). This is because our arguments, even after suitable modification, do not provide sufficient information about the density of units in the ring $\mathcal{O}_{{\mathbb{Q}}(\theta,\sqrt{-n})}$.

### v2

@@ 6. Further Remarks
Theorem 1.1 concerns itself with $26.5625\%$ of degree $3$ monic irreducible polynomials. This limitation arises because we only transferred lower bounds from $\mathcal{R}_{\theta,q}(\cdot,\cdot)$ to $B_{\mathbf{p}}(\cdot)$, where $q\in\{1,2\}$. It is important to note that sometimes $\mathcal{R}_{\theta,q}(\cdot,\cdot)$ can be empty, for example, if $q=1$ and $a_{1}a_{2}-a_{0}$ is even. To cover all cases of irreducible monic polynomials of degree $3$, we expect one would need to use sets $\mathcal{R}_{\theta,2^{t}}(\cdot,\cdot)$ for a suitable integer $t$ and employ enhanced parity arguments; however, we do not pursue these additional complications here. Furthermore, our arguments should generalise to study integer points on the Châtelet surface $y^{2}+nz^{2}=\mathbf{p}(x)$. Here our theory is stronger when $n\in\mathbb{N}$, as our unit arguments can be transferred quite readily, but when $n$ is negative, most of our theory breaks down (with the exception of an analogue for Lemma 1.3). This is because our arguments, even after suitable modification, do not provide sufficient information about the density of units in the ring $\mathcal{O}_{{\mathbb{Q}}(\theta,\sqrt{-n})}$.
@@ 7. Acknowledgements
8. Data availability statement
Data sharing is not applicable to this article as no datasets were generated or analysed during the current study.